Please answer this question........Oxidation state of Ni metal in ((Ni(NH3)6))(BF4)2 is

Grade:12

4 Answers

Arun
5 years ago
Oxidation no of Ni : +2
nomenclature of [Ni(NH3)6](BF4) 2 is “hexaamminenickel (||) tetrafluoroborate
Regards
Arun (askIITians forum expert)
Keerthana
5 years ago
Oxidation state is zero[Ni(NH3)6](BF4)2 in this let oxidation state of nickel be x thenX+6(0)-2(0)=0Since NH3 and BF4 have oxidation no (0)So x=0Oxidation state of Ni is zero
Keerthana
5 years ago
Oxidation state of Ni is 2Let oxidation no. Of Ni be xIt is a neutral compound so sum of oxidation states equal to 0X+6(0)+2(-1)=0 (since NH3 has oxi.no. 0 and BF4 has -1)So x=2Oxidation no. Of Ni is +2
Anjali kumari
5 years ago
Let oxidation state of Ni is x.Therefore x+(2×0)+2(-1) =0. or, x+0-2=0. or x=2. So oxidation state of Ni is +2.